OpenStudy (anonymous):

Can someone pls simplify √32+√18/3+√2 giving your answer in the form b√2 + c where b and c are intergers

OpenStudy (anonymous):

√32+√18/3+√2 = √(16*2)+√(9*2)/3+√2 =4√(2)+3√(2)/3+√2 =4√(2)+√(2)+√2 = 6√(2)+0 which is in the form of b√2 + c

OpenStudy (anonymous):

thanx